OpenStudy (anonymous):

find the equation of the line that has the given properties write equation in slope intercept form contains (3, 5) slope =7

OpenStudy (anonymous):

y=mx+b

OpenStudy (anonymous):

so 5=1(5)+7

OpenStudy (anonymous):

?

OpenStudy (anonymous):

x = 3, y = 5, and m = 7. You're solving for b. So, 5 = 7*3 + b.

OpenStudy (anonymous):

21+7=28

Directrix (directrix):

Use the point-slope general formula for a line: y - y1 = m(x - x1) where m is the slope of the line and (x1, y1) is the point on the line.

Directrix (directrix):

y - y1 = m(x - x1) y - (5) = 7( x - (3) ) y - 5 = 7x - 21 y = ? @romoor93

Directrix (directrix):

Let me know what you get, okay?

OpenStudy (anonymous):

y=7x-26

Directrix (directrix):

>y=7x-26 Almost ---- y - 5 = 7x - 21 y -5 + 5 = 7x - 21 + 5 y = 7x - ? --> What is -21 + 5?

OpenStudy (anonymous):

-16

Directrix (directrix):

Yes, so y = 7x - 16

OpenStudy (anonymous):

wow alot of steps but i think i get it

OpenStudy (anonymous):

thanks for steps

Directrix (directrix):

The steps are helpful to me. They will take you to the answer even if the walk is slow. :)